# https://re-ground.github.io

### v2

@@ Abstract
When an image generation process is guided by both a text prompt and spatial cues, such as a set of bounding boxes, do these elements work in harmony, or does one dominate the other? Our analysis of a pretrained image diffusion model that integrates gated self-attention into the U-Net reveals that spatial grounding often outweighs textual grounding due to the sequential flow from gated self-attention to cross-attention. We demonstrate that such bias can be significantly mitigated without sacrificing accuracy in either grounding by simply rewiring the network architecture, changing from sequential to parallel for gated self-attention and cross-attention. This surprisingly simple yet effective solution does not require any fine-tuning of the network but significantly reduces the trade-off between the two groundings. Our experiments demonstrate significant improvements from the original GLIGEN to the rewired version in the trade-off between textual grounding and spatial grounding.
@@
Figure 1: Comparison across Stable Diffusion (SD)[38], GLIGEN[26], and our ReGround. SD (2nd column) can generate an image aligned with the input prompt (shown below for each row), while it does not allow taking spatial constraints such as bounding boxes and labels. GLIGEN (3rd column) enables spatial grounding capability using gated self-attention, although it often disregards some descriptions in the input prompt due to bias towards bounding box information. Such trends also occur when only activating gated self-attention for 0.2 fraction of the initial denoising steps (4th column). Our ReGround (last column) resolves the issue of description omission while accurately reflecting the bounding box information.
@@ 1 Introduction
Among them, bounding boxes are extensively employed in various downstream image generation tasks [26, 51, 33, 29, 8, 4]. GLIGEN [26] is a pioneering work in terms of enhancing existing T2I models with the capability to incorporate additional spatial cues in the form of bounding boxes. Its core component, gated self-attention, is a simple attention module [47] that is plugged into each U-Net [39] layer of a pretrained T2I model, such as Stable Diffusion [38] and is trained to accurately position various entities in their designated areas. A notable advantage of GLIGEN is that the original parameters of the underlying model remain unchanged, inheriting the generative capability of the T2I model while introducing the novel functionality of spatial grounding using bounding boxes. This capability has been leveraged by numerous studies to facilitate high-quality, layout-guided image generation [10, 50, 33].
However, our analysis reveals that GLIGEN’s integration of the gated self-attention into an existing T2I model is not optimal for blending new spatial guidance from bounding boxes with the original textual guidance. It often leads to the omission of specific details from the text prompts. For instance, in the first row and the third column of Fig. 1, GLIGEN fails to reflect the description “giving a thumbs up” from the input text prompt. Also in the second row, a crucial detail in the text prompt, “draped with a colorful blanket”, is neglected in the output. We refer to this issue as description omission. Such outcomes imply that the current architectural design of GLIGEN does not effectively harmonize the new spatial guidance and the text conditioning in the given T2I model. Considering the widespread applications of GLIGEN in various layout-based generation tasks [10, 50, 33], these limitations represent a significant bottleneck.
@@
Building on this insight, we propose a straightforward yet impactful solution: network rewiring. This approach alters the relationship between the two grounding modules from sequential to parallel (Fig. 2-(c)). Remarkably, this network modification significantly reduces the grounding trade-off between textual and spatial groundings without necessitating any adjustments to the network parameters. Importantly, this rewiring does not require additional network training, extra parameters, or changes in computational load and time. Simply reconfiguring the attention modules of the pre-trained GLIGEN, originally trained with the sequential architecture, during inference dramatically enhances performance.
In our experiments with MS-COCO [28] and our newly introduced NSR-1K-GPT datasets, serving as ground truth, we demonstrate that rewiring the pretrained GLIGEN significantly reduces the trade-off between textual grounding and spatial grounding. This is evidenced by the evaluation of text prompt alignment (measured using CLIP Score [35] and PickScore [24]) and bounding box alignment (assessed by YOLO Score [49]). Furthermore, we show that our rewiring also leads to better outcomes in other frameworks using GLIGEN as a backbone, including BoxDiff [50].
@@ 2.1 Zero-Shot Guidance in Diffusion Models
The progress in diffusion models [15, 42, 41] has significantly elevated the capabilities of text-to-image (T2I) generation, resulting in foundation models [38, 34, 37, 36, 6] that exhibit remarkable generative performance. Leveraging the robust performance of these models, recent studies [26, 51, 50, 33, 29, 8, 4, 3, 23, 5, 9, 53] have introduced efficient guidance techniques designed to specifically improve the image generation process. Notably, numerous works [26, 51, 50, 33, 29, 8, 40] focus on the internal architecture (Fig. 2-(a)) of the denoising U-Net of Latent Diffusion Models (LDMs) [38], where self-attention and cross-attention modules are intertwined to facilitate inter-pixel communication and text conditioning. The self-attention of U-Net can be utilized to improve image quality [17], facilitate image translation [46], and support image editing tasks [7]. Since the text condition is integrated into cross-attention, the intermediate attention map has also been widely leveraged to improve text faithfulness [11] or enable spatial manipulation of the generation process [32] using text prompts. Recently, FreeU [40] analyzed the contributions of the backbone and residuals of the U-Net and proposed a free-lunch strategy to enhance image quality: reweighting the backbone and residual features maps. In contrast to previous works that only deal with self- and cross-attention in standard LDMs, we introduce a method to enhance GLIGEN [26] by reconnecting its gated self-attention with the other attention modules, thereby achieving performance improvement in zero-shot without any tuning of the network parameters.
@@ 2.2 Layout-Guided Image Generation
The use of layouts, particularly in the form of bounding boxes, has become a popular intermediary to bridge the gap between textual inputs and the images generated [54, 43, 25, 27, 16, 19, 44, 12, 52]. Layout2Im [54] learns to sample object latent codes from a normal distribution, eliminating the need for predicting instance masks as done in prior works [16, 19]. LostGAN [43] improves the ability to control the style of each object by devising an extension of the feature normalization layer used in StyleGAN [21, 22, 20], while OC-GAN  [44] incorporates the spatial relationships between objects using a scene-graph representation. LAMA [27] introduces a mask adaptation module that mitigates the semantic ambiguity arising from overlaps in the input layout. While these developments have greatly improved user control over image generation, their applicability is confined to the categories found in the training data, such as those of the MS-COCO [28] dataset.
In contrast, recent studies [26, 51, 8, 33, 50, 56, 5] have extended layout-guided image generation towards open-vocabulary, building on the advancements of foundational text-to-image (T2I) models [38]. Training-free approaches [50, 3, 33, 8, 5] aim to improve the spatial grounding of T2I models through straightforward guidance mechanisms. GLIGEN [26], on the other hand, introduces gated self-attention, which is injected into the U-Net architecture of the Latent Diffusion Model [38], and is trained to equip the underlying model with spatial grounding ability. Given the simple architecture of GLIGEN and its robust grounding accuracy with the input bounding boxes, numerous studies [33, 50, 55] build upon its framework and propose further refinements to increase performance. In this work, we identify and address a significant performance bottleneck in GLIGEN related to description omission and propose a simple yet effective solution.
@@ 3 Background — Latent Diffusion Models [38]
Rombach et al. [38] proposed Latent Diffusion Models (LDM), a text-to-image (T2I) diffusion model using a U-Net as the noise prediction network. It is trained to generate an image from an input text prompt by predicting the noise $\epsilon(\mathbf{x}_{t},t,c)$ conditioned both on the timestep $t$ and the text embedding $c$, which is obtained by encoding a text prompt using a pretrained text encoder [35, 18]. Each layer of LDM’s U-Net consists of three core components: a convolutional residual block, followed by a self-attention (SA), and a cross-attention (CA) module (Fig. 2-(a)). In each $l$-th layer of the U-Net, its residual block first extracts intermediate visual features $F=(f_{1},...,f_{N_{l}})^{T}$ from the output of the previous layer. Following this, a self-attention module facilitates interaction between the features in $F$. Subsequently, a cross-attention module enables the interaction between each visual feature $f_{i}$ and the text embedding $c$. Throughout this process, the output feature of the previous module is also forwarded through a residual connection, as illustrated in lines 4-5 and 7 of Alg. 1 and also in Fig. 2-(a).
@@ 4.1 Gated Self-Attention
Li et al.[26] propose a plug-in spatial grounding module, named gated self-attention, which adopts the gated attention mechanism[1] to equip a pretrained T2I diffusion model [38] with the spatial grounding ability (Fig. 2-(b)). Given a set of bounding boxes in the image space and text labels for each of them, let $b_{i}$ be the $xy$-coordinates of the $i$-th bounding box’s top-left and bottom-right corners, and $p_{i}$ be the corresponding text label. Then, the $i$-th grounding token is defined as $g_{i}:=\mathcal{G}\left(\mathcal{T}\left(p_{i}\right),\mathcal{F}\left(b_{i}\right)\right)$, where $\mathcal{T}(\cdot)$ is a pretrained text encoder [35, 18], $\mathcal{F}(\cdot)$ is the Fourier embedding [30, 45] and $\mathcal{G}(\cdot,\cdot)$ is a shallow MLP network that concatenates the two given embeddings, respectively. Given a set of grounding tokens $\{g_{i}\}$, gated self-attention learns the self-attention among the unified feature set $(f_{1},...,f_{N_{l}},g_{1},...,g_{M})$ where $\{f_{i}\}$ is the set of intermediate visual features in the $l$-th layer of U-Net, and $M$ is the number of bounding boxes.
As shown in Fig. 2-(b), gated self-attention receives the output of the self-attention as its input and forwards the output feature to the cross-attention module. By incorporating gated self-attention into each layer of the U-Net, the model enables the placement of the entity specified in the text label $p_{i}$ at the location indicated by the bounding box $b_{i}$. Note that the integration of gated self-attention does not require training the network from scratch or fine-tuning it, but can be accomplished simply by training the gated self-attention parameters while keeping all other parameters in the backbone model frozen.
@@ 7       $F\leftarrow\mathbf{x}_{t}$
             $F_{\text{RN}}\leftarrow\text{Conv}(F)+F$;
@@
             $F_{\text{SA}}\leftarrow\text{SA}(F_{\text{RN}})+F_{\text{RN}}$;
@@ 4.2 Description Omission
Despite its high accuracy in spatial grounding, GLIGEN [26] frequently struggles to capture essential attributes specified in the input text prompt. As illustrated in Fig. 3, the image on the left shows a person and a skateboard accurately placed in their designated regions. However, a critical detail from the input text prompt, “black and white photography”, is conspicuously absent in the output image. This discrepancy often emerges when the input comprises distinct but equally important descriptions regarding the image, presented through text prompts and bounding boxes. Such omissions not only fail to convey the stylistic intent of the image but also tend to overlook significant objects mentioned within the text prompt. Additional examples of this problem are showcased in Fig. 1, where, for example, the second row demonstrates the absence of a “blanket” in the generated image, a key element from the text prompt. This limitation significantly hampers GLIGEN’s fidelity to user-provided text prompts, a challenge we term as description omission.
@@ 5 ReGround: Rewiring Attention Modules
Figure 3: (a) Images generated by GLIGEN [26] while varying the activation duration of gated self-attention $\gamma$ in scheduled sampling (Sec. 5.1). The red words in the text prompt denote the words used as labels of the input bounding boxes. Note that to reflect the underlined description in the text prompt in the final image, $\gamma$ must be decreased to 0.1, which compromises spatial grounding accuracy. (b) In contrast, our ReGround reflects the underlined phrase even when $\gamma=1.0$, therefore achieving high accuracy in both textual and spatial grounding.
@@ 5.1 Impact of Gated Self-Attention on Textual Grounding
As the issue of description omission arises due to the newly added gated self-attention in GLIGEN [26], we first attempt to mitigate the impact of gated self-attention by using scheduled sampling [26], activating gated self-attention only in a few initial steps of the denoising process. This approach is inspired by the observation that the coarse structure of the final image is established within the first few time steps. The scheduling is applied by setting the weight of gated self-attention $\beta_{t}$ (line 6 of Alg. 1) as
@@
Fig. 3-(a) shows an example of generated images while incrementally adjusting $\gamma$ from 1.0 to 0.0. As $\gamma$ is reduced from 1.0 to 0.0, the details specified in the text prompt, “a black and white photograph”, begin to be reflected starting at $\gamma=0.1$, demonstrating that longer activation of gated self-attention may interfere with the alignment of the output image with the text prompt. However, as gated self-attention is activated for shorter durations, the spatial grounding diminishes, as shown in the objects’ reduced alignment with the input bounding box. This phenomenon illustrates the inherent trade-off between spatial and textual grounding, which cannot be resolved by controlling the duration of gated self-attention activation.
@@ 5.2 Impact of Cross-Attention on Spatial Grounding
We also investigate whether cross-attention has influence on the spatial grounding. To examine it, we conduct a toy experiment by removing cross-attention modules in GLIGEN [26], so that the output of gated self-attention is directly passed onto the next layer of the U-Net. This is equivalent to changing line 7 of Alg. 1 to $F\leftarrow F_{GSA}$.

### v3

# ReGround: Improving Textual and Spatial Grounding at No Cost

@@ Abstract
When an image generation process is guided by both a text prompt and spatial cues, such as a set of bounding boxes, do these elements work in harmony, or does one dominate the other? Our analysis of a pretrained image diffusion model that integrates gated self-attention into the U-Net reveals that spatial grounding often outweighs textual grounding due to the sequential flow from gated self-attention to cross-attention. We demonstrate that such bias can be significantly mitigated without sacrificing accuracy in either grounding by simply rewiring the network architecture, changing from sequential to parallel for gated self-attention and cross-attention. This surprisingly simple yet effective solution does not require any fine-tuning of the network but substantially reduces the trade-off between the two groundings. Our experiments demonstrate significant improvements from the original GLIGEN to the rewired version in the trade-off between textual grounding and spatial grounding. The project webpage is at https://re-ground.github.io.
@@
Figure 1: Comparison across Stable Diffusion (SD)[42], GLIGEN[28], and our ReGround. SD (2nd column) can generate an image aligned with the input prompt (shown below each row), while it does not allow taking spatial constraints such as bounding boxes and labels. GLIGEN (3rd column) enables spatial grounding using gated self-attention, although it often disregards some descriptions in the input prompt due to a bias towards bounding box conditions. Such trends also occur when only activating gated self-attention for 0.2 fraction of the initial denoising steps (4th column). Our ReGround (last column) resolves the issue of description omission while accurately reflecting the bounding box information.
@@ 1 Introduction
Among them, bounding boxes are extensively employed in downstream image generation tasks [28, 56, 37, 33, 8, 4]. GLIGEN [28] is a pioneering work in terms of enhancing existing T2I models with the capability to incorporate additional spatial cues in the form of bounding boxes. Its core component, gated self-attention, is a simple attention module [51] that is plugged into each U-Net [43] layer of a pretrained T2I model such as Stable Diffusion [42], and is trained to accurately position various entities in their designated areas. A notable advantage of GLIGEN is that the original parameters of the underlying model remain unchanged, inheriting the generative capability of the T2I model while introducing the novel functionality of spatial grounding using bounding boxes. This capability has been leveraged by numerous studies to facilitate high-quality, layout-guided image generation [12, 55, 9, 32].
However, our analysis reveals that GLIGEN’s integration of the gated self-attention into an existing T2I model is not optimal for blending new spatial guidance from bounding boxes with the original textual guidance. It often leads to the omission of specific details from the text prompts. For instance, in the first row and third column of Fig. 1, GLIGEN fails to reflect the description “low poly illustration” from the input text prompt. Also in the second row, a crucial detail in the text prompt, “draped with a colorful blanket”, is neglected in the output image. We refer to this issue as description omission. Such outcomes imply that the current architectural design of GLIGEN does not effectively harmonize the new spatial guidance and the text conditioning in the given T2I model. Considering the widespread applications of GLIGEN in various layout-based generation tasks [12, 55, 37, 61, 54, 9, 32], these limitations represent a significant bottleneck.
@@
Building on this insight, we propose a straightforward yet impactful solution: network rewiring. This approach alters the relationship between the two grounding modules from sequential to parallel (Fig. 2-(c)). Remarkably, this network modification significantly reduces the grounding trade-off between textual and spatial groundings without necessitating any adjustments to the network parameters. Importantly, this rewiring does not require additional network training, extra parameters, or changes in computational load and time. Simply reconfiguring the attention modules of the pretrained GLIGEN, originally trained with the sequential architecture, during inference dramatically enhances performance.
In our experiments on MS-COCO [31] and our newly introduced NSR-1K-GPT datasets, we demonstrate that rewiring the pretrained GLIGEN substantially reduces the trade-off between textual and spatial groundings. This is evidenced by the evaluation of text prompt alignment (measured using CLIP score [39], PickScore [26] and user study) and bounding box alignment (assessed by YOLO score [53]). Furthermore, we show that our rewiring also leads to better outcomes in other frameworks using GLIGEN as a backbone, including BoxDiff [55].
@@ 2.1 Zero-Shot Guidance in Diffusion Models
The progress in diffusion models [17, 46, 45] has significantly elevated the capabilities of text-to-image (T2I) generation, resulting in foundation models [42, 38, 41, 40, 6] that exhibit remarkable generative performance. Leveraging the robust performance of these models, recent studies [28, 56, 55, 37, 33, 8, 4, 3, 25, 5, 11, 58] have introduced efficient guidance techniques designed to further improve the image generation process. Notably, numerous works [28, 56, 55, 37, 33, 8, 44] focus on the internal architecture (Fig. 2-(a)) of the denoising U-Net of Latent Diffusion Models (LDMs) [42], where self-attention and cross-attention modules are intertwined to facilitate inter-pixel communication and text conditioning. The self-attention of U-Net can be utilized to improve image quality [19] or facilitate image translation [50] and image editing tasks [7]. Since text conditions are integrated via cross-attention, the intermediate attention maps have been leveraged to improve text faithfulness [13] or enable spatial manipulation of the generation process [36]. Recently, FreeU [44] analyzed the contributions of the backbone and residuals of the U-Net and proposed a free-lunch strategy to enhance image quality: reweighting the backbone and residual features maps. In contrast to previous works that only deal with self- and cross-attention in standard LDMs, we introduce a method to enhance GLIGEN [28] by reconnecting its gated self-attention with the other attention modules, thereby achieving performance improvement in zero-shot without any tuning of the network parameters.
@@ 2.2 Layout-Guided Image Generation
The use of layouts, particularly in the form of bounding boxes, has become a popular intermediary to bridge the gap between textual inputs and the images generated [60, 47, 27, 29, 18, 21, 48, 14, 57]. Layout2Im [60] samples object latent codes from a normal distribution, eliminating the need to predict instance masks as done in prior works [18, 21]. LostGAN [47] controls the style of each object by devising an extension of the feature normalization layer used in StyleGAN [23, 24, 22], while OC-GAN  [48] incorporates the spatial relationships between objects using a scene-graph representation. LAMA [29] introduces a mask adaptation module that mitigates the semantic ambiguity arising from overlaps in the input layout. While these developments have greatly improved user control over image generation, their applicability is confined to the categories found in the training data, such as those of the MS-COCO [31] dataset.
In contrast, recent studies [28, 56, 8, 37, 55, 62, 5, 10] have extended layout-guided image generation towards open-vocabulary, building on the advancements of foundational text-to-image (T2I) models [42]. Training-free approaches [55, 3, 37, 8, 5] aim to improve the spatial grounding of T2I models through straightforward guidance mechanisms. GLIGEN [28], on the other hand, introduces gated self-attention, which is injected into the U-Net architecture of the Latent Diffusion Model [42], and is trained to equip the underlying model with spatial grounding abilities. Given the simple architecture of GLIGEN and its robust grounding accuracy with the input bounding boxes, numerous studies [37, 55, 61, 54] build upon its framework and propose further refinements to increase performance. In this work, we identify and address a significant performance bottleneck in GLIGEN related to description omission and propose a simple yet effective solution.
@@ 3 Background — Latent Diffusion Models [42]
Rombach et al. [42] proposed Latent Diffusion Model (LDM), a text-to-image (T2I) diffusion model with a U-Net as the noise prediction network. It is trained to generate an image from an input text prompt by predicting the noise $\epsilon(\mathbf{x}_{t},t,c)$ conditioned both on the timestep $t$ and the text embedding $c$. Each layer of LDM’s U-Net consists of three core components: a convolutional residual block, followed by a self-attention (SA), and a cross-attention (CA) module (Fig. 2-(a)). In each $l$-th layer of the U-Net, its residual block first extracts intermediate visual features $F=(f_{1},...,f_{N_{l}})^{T}$ from the output of the previous layer. The self-attention module then facilitates interaction between the features in $F$. Subsequently, the cross-attention module enables the interaction between each visual feature $f_{i}$ and the text embedding $c$. Throughout this process, the output feature of the previous module is also forwarded through a residual connection, as illustrated in lines 4-5 and 7 of Alg. 1 and also in Fig. 2-(a).
@@ 4.1 Gated Self-Attention
Li et al.[28] propose a plug-in spatial grounding module, named gated self-attention, which adopts the gated attention mechanism[1] to equip a pretrained T2I diffusion model [42] with spatial grounding capabilities (Fig. 2-(b)). Given a set of bounding boxes and text labels for each of them, let $b_{i}$ be the $xy$-coordinates of the $i$-th bounding box’s top-left and bottom-right corners, and $p_{i}$ be the corresponding text label. Then, the $i$-th grounding token is defined as $g_{i}:=\mathcal{G}\left(\mathcal{T}\left(p_{i}\right),\mathcal{F}\left(b_{i}\right)\right)$, where $\mathcal{T}(\cdot)$ is a pretrained text encoder [39, 20], $\mathcal{F}(\cdot)$ is the Fourier embedding [34, 49] and $\mathcal{G}(\cdot,\cdot)$ is a shallow MLP network that concatenates the two given embeddings, respectively. Given a set of grounding tokens $\{g_{i}\}$, gated self-attention learns the self-attention among the unified feature set $(f_{1},...,f_{N_{l}},g_{1},...,g_{M})$ where $\{f_{i}\}$ is the set of intermediate visual features in the $l$-th layer of U-Net, and $M$ is the number of bounding boxes.
As shown in Fig. 2-(b), gated self-attention receives the output of the self-attention along with the residual features as its input and forwards the output features to the cross-attention module. By incorporating gated self-attention into each layer of the U-Net, the model enables the placement of the entity specified in the text label $p_{i}$ at the location indicated by the bounding box $b_{i}$. Note that the integration of gated self-attention does not require training the network from scratch or fine-tuning it, but can be accomplished simply by training the gated self-attention parameters while keeping all other parameters in the backbone model frozen.
@@ 7       $F\leftarrow\mathbf{x}_{t}$
             $F_{\text{RS}}\leftarrow\text{Conv}(F)+F$;
@@
             $F_{\text{SA}}\leftarrow\text{SA}(F_{\text{RS}})+F_{\text{RS}}$;
@@ 4.2 Description Omission
Despite its high accuracy in spatial grounding, GLIGEN [28] frequently struggles to capture essential attributes specified in the input text prompt. As illustrated in Fig. 3, the leftmost image shows “a person” and “a skateboard” accurately placed in their designated regions. However, a critical detail from the input text prompt, “black and white photography”, is absent in the output image. This discrepancy often emerges when the input comprises distinct but equally important descriptions regarding the image, presented through text prompts and bounding boxes. Such omissions not only fail to convey the stylistic intent of the image but also tend to overlook significant objects mentioned within the text prompt. Additional examples of this problem are showcased in Fig. 1, where the second row demonstrates the absence of a “blanket” in the generated image, a key element from the text prompt. This limitation significantly hampers GLIGEN’s fidelity to user-provided text prompts, a challenge we term as description omission.
@@ 5 ReGround: Rewiring Attention Modules
Figure 3: (a) Images generated by GLIGEN [28] with varying activation duration of gated self-attention $\gamma$ in scheduled sampling (Sec. 5.1). The red words in the text prompt denote the words used as labels of the input bounding boxes. Note that for GLIGEN to reflect the underlined description in the text prompt in the final image, $\gamma$ must be decreased to 0.1, which compromises spatial grounding accuracy. (b) In contrast, our ReGround reflects the underlined phrase even when $\gamma=1.0$, therefore achieving high accuracy in both textual and spatial grounding.
@@ 5.1 Impact of Gated Self-Attention on Textual Grounding
As the issue of description omission arises due to the newly added gated self-attention in GLIGEN [28], we first attempt to mitigate the impact of gated self-attention by using scheduled sampling [28], activating gated self-attention only in a few initial steps of the denoising process. This approach is inspired by the observation that the coarse structure of the final image is established within the first few denoising steps. The scheduling is applied by setting the weight of gated self-attention $\beta_{t}$ (line 6 of Alg. 1) as
@@
Fig. 3-(a) shows an example of generated images while incrementally adjusting $\gamma$ from 1.0 to 0.0. As $\gamma$ is reduced from 1.0 to 0.0, the details specified in the text prompt, “a black and white photograph”, begin to be reflected starting at $\gamma=0.1$, demonstrating that longer activation of gated self-attention may interfere with the alignment of the output image with the text prompt. However, as gated self-attention is activated for shorter durations, the spatial grounding diminishes, as shown in the objects’ reduced alignment with the input bounding boxes. This phenomenon illustrates the inherent trade-off between spatial and textual grounding, which cannot be resolved by controlling the duration of gated self-attention activation.
@@ 5.2 Impact of Cross-Attention on Spatial Grounding
We also investigate whether cross-attention has influence on spatial grounding. For this, we conduct a toy experiment by removing cross-attention modules in GLIGEN [28], allowing the output of the gated self-attention to be directly passed to the next layer of the U-Net. This modification is equivalent to changing line 7 of Alg. 1 to $F\leftarrow F_{GSA}$.